if nikos and noel live 39 miles apart and noel rides his bike at a rate of 14 mph and nikos rides his at a rate of 12 mph and they both start biking towards eachother at 3p.m at what time will they meet?
Answer by checkley77(12844)

D=RT
39=(14+12)T
39=26T
T=39/26
T=1.5 HOURS.
3+1.5=4:30 THEY WILL MEET.